OpenStudy (anonymous):

A model car has a scale of 1:25. The length of the model when built is 9.6 in. What is the length of the actual car? a) 20 feet b) 20 inches c)19.2 feet d) 25 feet

OpenStudy (johnweldon1993):

Well it is important to realize...for every inch of the model car, there is 25 inches in the real car. So if the model is 9.6 inches...the real car is \(\large 9.6 \times 25 = 240 in\) Now we obviously dont have that as an answer...however if we convert the 240 inches to feet...I believe you will see the correct answer

OpenStudy (triciaal):

set up a ratio 1/25 = 9.6/x where x is the actual solve for x